OpenStudy (anonymous):

Simplify the expression below. If your answer is not a whole number, enter it as a fraction in lowest terms, using the slash mark ( / ) as the fraction bar. (17^1/8)^8

OpenStudy (anonymous):

PEMDAS (Parentheses, Exponents, Multiplication, Division, Addition, Subtraction) That is the order of operations

OpenStudy (marihelenh):

\[(17^{\frac{ 1 }{ 8 }})^{8}\]

OpenStudy (tkhunny):

\(\left(a^{b}\right)^c = a^{bc}\)

OpenStudy (anonymous):

Remember, when it is an exponent to the power of an exponent, you multiply them.

OpenStudy (anonymous):

ex: (2^2)^8 = 2^16

OpenStudy (anonymous):

so, what is 1/8 x 8? that will be the power of 17. Then simplify.

OpenStudy (anonymous):

1/8 times 8 is 1 @robjoe

OpenStudy (anonymous):

so it will be 17 ?

OpenStudy (anonymous):

@norajayy that is right!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem!

OpenStudy (anonymous):

would i do the same thing for (6^7)^1/7 7 times 1/7 ? @robjoe

OpenStudy (anonymous):

the answer will be six

OpenStudy (anonymous):

@norajayy yes